# Caption This, Reason That: VLMs Caught in the Middle

### v1

@@ A.5 Captioning Method Validation
Task Qwen 72B-Base Qwen 72B-SC Qwen 72B-PC 4o-Base 4o-SC 4o PC
Perception (Cat) 89.00±4.36 -2.67±3.89 +2.00±4.00 90.33±3.36 -2.00±3.64 +0.67±3.26
Perception (Loc) 70.50±6.27 +9.17±4.54 +29.50±0.94 89.00±3.55 +8.33±1.91 +11.00±0.63
Feature Attention 80.44±2.59 +8.15±1.70 +19.22±0.43 87.19±1.78 +8.89±1.04 +12.81±0.14
Spatial Attention 80.56±2.58 +5.44±1.85 +15.44±1.29 75.93±2.28 +15.33±1.51 +19.85±1.08
Memory (Cat) 87.69±1.29 -1.42±1.10 +6.59±0.91 91.47±0.89 +1.73±0.81 +3.44±0.70
Memory (Loc) 58.44±1.97 +17.78±1.39 +36.47±0.88 83.47±1.21 +9.89±0.81 +12.22±0.66
CVR-Cat-L 80.00±7.77 -4.67±6.84 +20.00±1.85 91.33±4.56 +2.00±4.09 +8.67±1.25
CVR-Loc-L 52.00±9.61 +30.67±6.04 +47.00±2.64 66.00±7.50 +27.33±4.09 +34.00±1.25
CVR-Cat-M 88.00±6.41 -11.33±6.72 +6.00±4.85 96.67±3.07 -7.33±4.98 +0.67±2.81
CVR-Loc-M 48.00±9.61 +28.67±6.72 +42.00±5.96 82.67±6.04 +7.33±4.84 +14.67±2.81
CVR-Cat-H 69.83±6.33 -7.83±5.46 +8.17±5.71 83.67±4.18 -2.00±4.37 +2.67±3.89
CVR-Loc-H 40.97±6.74 +24.36±5.35 +50.03±4.00 64.67±5.38 +18.67±4.21 +30.33±2.52
A.6 PAM & CVR vs MMMU-Pro Performance
A.7 Prompt & Script Examples
{minted}
[fontsize=, linenos=false]json "messages": [ "role": "user", "content": "In this task, we will show you a series of frame images. Each frame will either be blank (delay frame) or contain one or more 3D objects. The objects will always be from one of eight categories: benches, boats, cars, chairs, couches, lighting, planes, and tables. For each category, there are eight unique objects that could be used in the task. Any object sampled will be displayed as an image taken from a random viewing angle. The objects will be placed in one of four locations: top left, top right, bottom left, and bottom right. If there are multiple objects on a single frame, only one of them would be specified in the task instruction by either its location or its category. A written instruction will be provided. Your goal is to follow the instructions and answer the question contained in the instructions. Answers will always be one of the following: true, false, bottom right, bottom left, top left, top right, benches, boats, cars, chairs, couches, lighting, planes, tables.
Please solve the following task: Task instruction: observe object 1 in frame 1, observe object 2 in frame 2, observe object 3 in frame 3, observe object 4 in frame 4, observe object 5 in frame 5, delay, observe object 6 in frame 7, delay, observe object 7 in frame 9, if identity of object 3 equals identity of object 2, then location of object 7 not equals location of object 6 and identity of object 5 equals identity of object 4? else location of object 1?
Here are the corresponding frames: <image><image><image><image><image><image><image> <image><image>
What is the correct answer to this task? (bottom right, bottom left, top left, top right). Provide your answer here: " , "role": "assistant", "content": "top right" ], "images": [ "../trial0/frames/epoch0.png", "../trial0/frames/epoch0.png", "../trial0/frames/epoch0.png", "../trial0/frames/epoch0.png", "../trial0/frames/epoch0.png", "../trial0/frames/epoch0.png", "../trial0/frames/epoch0.png", "../trial0/frames/epoch0.png", "../trial0/frames/epoch0.png" ]
@@ A.7 Prompt & Script Examples
{minted}
[fontsize=]text "Frame 1: A chairs located at the top left", "Frame 2: A chairs located at the top right", "Frame 3: A benches located at the top left", "Frame 4: A benches located at the bottom right", "Frame 5: A boats located at the bottom left", "Frame 6: A benches located at the bottom right", "Frame 7: delay frame", "Frame 8: delay frame", "Frame 9: A planes located at the top left"
@@
{minted}
[fontsize=, linenos=false,breaklines]python def caption_image(image_path, model, config): # Prepare a prompt asking the model to caption the image: caption_prompt = [ "role": "user", "content":[ "type": "text", "text": "Please provide a concise caption for the given image,  including what the location of each the object in the images  are and what the category of each object is. Each image either  is blank (a delay frame) or contains one or more 3D objects from  one of eight categories: benches, boats, cars, chairs, couches,  lighting, planes, and tables. The object is placed in one of four  locations: top left, top right, bottom left, or bottom right." , "type": "image_url", "image_url": "url": f"data:image/png;base64,encode_image(image_path)" ]
# Call the model to get a caption response = model.chat.completions.create( model=config.get("oai-model", "gpt-4o-mini"), messages=caption_prompt, max_tokens=1024 )
# The model’s caption is expected in the response. caption = response.choices[0].message.content.strip() return caption
@@
{minted}
[fontsize=, linenos=false, breaklines]python def evaluate_model(task_base, trial_num, model, config): trial_base = f"task_base/trialtrial_num/frames/" task_info_path = os.path.join(trial_base, "new_task_info.json")
with open(task_info_path, ’r’) as file: data = json.load(file)
# Extract the instruction and answers instruction = data[’new_instruction’] answers = data[’answers’]
# Find the images pattern = ’epoch*.png’ matching_files = [f for f in os.listdir(trial_base) if fnmatch.fnmatch(f, pattern)] matching_files.sort()
images = [os.path.join(trial_base, file) for file in matching_files]
# —————————————————– # NEW PART: First, caption each image # —————————————————– captions = [] for image_path in images: caption = caption_image(image_path, model, config) captions.append(caption)
# Now we have a list of captions, one per frame # Incorporate these captions into the final prompt # Instead of sending images to the model in the final task, we send their captions:
# Prepare the final evaluation prompts # The prompt should now describe that we have textual descriptions of each frame: possible_answers = reduce(lambda x, y: f"x, y", get_all_possible_ans(answers[-1]))
prompts = [ f"""In this task, we will show you a series of frames described by captions instead of direct images. Each frame either is blank (a delay frame) or contains a 3D object from one of eight categories: benches, boats, cars, chairs, couches, lighting, planes, and tables. The object is placed in one of four locations: top left, top right, bottom left, or bottom right.
A written instruction is provided, and you must follow that instruction to find the correct answer. The valid answers are always one of the following: true, false, bottom right, bottom left, top left, top right, benches, boats, cars, chairs, couches, lighting, planes, tables.
Task instruction: instruction
@@
{minted}
[fontsize=, linenos=false]python … messages = [ "role": "user", "content": [ "type": "text", "text": prompts[0], , "type": "text", "text": prompts[1], , ] ]
print(messages) instruction = instruction.join([f"Frame i+1: cap" for i, cap in enumerate(captions)])
# Get the response for the final evaluation response = model.chat.completions.create( model=config.get("oai-model", "gpt-4o-mini"), messages=messages, max_tokens=config.get("max_new_tokens", 50), )
return instruction, answers[-1], response.choices[0].message.content

### v2

@@ A.5 Captioning Method Validation
Table A.5.1: PAM and CVR performance of Qwen2.5-VL-72B using different captioning methods. The Base column shows absolute percentage accuracies ($\pm$ standard deviation). Subsequent columns show the percentage change from the Base method. Positive changes are shown in green, negative in red. SC: Self-captioning, PC: Pre-captioned. Asterisks denote statistical significance: * p < 0.05, ** p < 0.01, *** p < 0.001.
Task Qwen 72B Base Qwen SC Qwen PC
Perception (Cat) 89.00±4.36 -2.67±3.89 +2.00±4.00
Perception (Loc) 70.50±6.27 +9.17±4.54 +29.50±0.94
Feature Attention 80.44±2.59 +8.15**±1.70 +19.22***±0.43
Spatial Attention 80.56±2.58 +5.44**±1.85 +15.44***±1.29
Memory (Cat) 87.69±1.29 -1.42*±1.10 +6.59***±0.91
Memory (Loc) 58.44±1.97 +17.78***±1.39 +36.47***±0.88
CVR-Cat-L 80.00±7.77 -4.67±6.84 +20.00±1.85
CVR-Loc-L 52.00±9.61 +30.67±6.04 +47.00±2.64
CVR-Cat-M 88.00±6.41 -11.33±6.72 +6.00±4.85
CVR-Loc-M 48.00±9.61 +28.67±6.72 +42.00±5.96
CVR-Cat-H 69.83±6.33 -7.83±5.46 +8.17±5.71
CVR-Loc-H 40.97±6.74 +24.36±5.35 +50.03±4.00
Table A.5.2: PAM and CVR performance of GPT-4o using different captioning methods. The Base column shows absolute percentage accuracies ($\pm$ standard deviation). Subsequent columns show the percentage change from the Base method. Positive changes are shown in green, negative in red. SC: Self-captioning, PC: Pre-captioned. Asterisks denote statistical significance: * p < 0.05, ** p < 0.01, *** p < 0.001.
Task 4o Base 4o SC 4o PC
Perception (Cat) 90.33±3.36 -2.00±3.64 +0.67±3.26
Perception (Loc) 89.00±3.55 +8.33±1.91 +11.00*±0.63
Feature Attention 87.19±1.78 +8.89***±1.04 +12.81***±0.14
Spatial Attention 75.93±2.28 +15.33***±1.51 +19.85***±1.08
Memory (Cat) 91.47±0.89 +1.73±0.81 +3.44***±0.70
Memory (Loc) 83.47±1.21 +9.89***±0.81 +12.22***±0.66
CVR-Cat-L 91.33±4.56 +2.00±4.09 +8.67±1.25
CVR-Loc-L 66.00±7.50 +27.33±4.09 +34.00±1.25
CVR-Cat-M 96.67±3.07 -7.33±4.98 +0.67±2.81
CVR-Loc-M 82.67±6.04 +7.33±4.84 +14.67±2.81
CVR-Cat-H 83.67±4.18 -2.00±4.37 +2.67±3.89
CVR-Loc-H 64.67±5.38 +18.67*±4.21 +30.33*±2.52
A.6 Additional Fine-tuning Results
Table A.6.1: PAM and CVR performance of Qwen2.5-VL-32B base model and fine-tuned variants using LoRA with different amounts of training data. The Base column shows absolute percentage accuracies ($\pm$ standard deviation). Subsequent columns show the percentage change in accuracy from the Base method ($\pm$ standard deviation). Positive changes are shown in green, and negative in red. Asterisks denote statistical significance: * p < 0.05, ** p < 0.01, *** p < 0.001.
Task 32B Base 32B LoRA 1k 32B LoRA 100k
Memory (Cat) 85.38±1.13 -1.91±1.19 +2.61***±1.04
Memory (Loc) 48.89±1.63 +8.44***±1.61 +16.80***±1.55
Perception (Cat) 87.67±3.73 -1.33±3.89 -1.00±3.85
Perception (Loc) 49.67±5.62 +19.33±5.21 +29.67*±4.57
Feature Attention 76.13±2.27 -11.13*±2.54 -10.31*±2.53
Spatial Attention 80.98±2.09 -0.54±2.12 +1.52±2.03
CVR-Cat-H 59.39±5.52 -6.06±5.61 +16.94*±4.79
CVR-Loc-H 35.33±5.38 +5.33±5.52 +26.33*±5.47
CVR-Cat-M 72.67±7.07 -19.33±7.88 -4.67±7.38
CVR-Loc-M 46.00±7.88 +6.00±7.89 +15.33±7.70
CVR-Cat-L 83.33±5.95 -24.67±7.78 -18.00±7.53
CVR-Loc-L 48.67±7.90 +10.00±7.78 +15.33±7.59
A.7 PAM & CVR vs MMMU-Pro Performance
A.8 PAM & CVR vs VQAv2 Performance
Figure A.8.1: Scatter plots comparing average PAM task (Perception, Attention, Memory) performance and CVR task performance against VQAv2 performance across different variants of Qwen2.5 VL models. Each point represents a different configuration. The x-axis shows the average accuracy on the specified PAM task category (averaging Loc and Cat), and the y-axis shows the overall VQAv2 accuracy.
A.9 Attention Image Capacity Analysis
To investigate potential attention capacity limitations introduced by the presence of images during reasoning, we compared the average attention allocated to ground-truth caption tokens across the chain-of-thought (CoT) tokens generated by Qwen2.5-VL-7B in both PC (pre-captioned) and PC-I (pre-captioned with images) trials. Appendix Figure A.9.1 shows the inclusion of images significantly reduces the attention a given ground-truth caption token receives while the model outputs a CoT. These findings may reflect a potential attention capacity issue imposed by images during reasoning, which improvements during CoT training could remedy.
Figure A.9.1: Comparison of average Qwen2.5-VL-7b attention scores for ground-truth caption tokens across CoT tokens during PC and PC-I prompts of PAM tasks.
A.10 Inference Runtime Analysis
To assess the computational overhead of our proposed Self-Captioning (SC) method, we measured the average inference runtime and compared it against the baseline direct-input method. We conducted these analyses on a Qwen2.5-VL-7B model running on a single NVIDIA A5000 GPU. We timed 10 trials for each task. The results, detailed in Table A.10.1, show that the SC method incurs a significant runtime cost, taking approximately twice as long as the base method for both simple and complex tasks.
Table A.10.1: Inference runtime comparison for the Self-Captioning (SC) vs. Base method on Qwen2.5-VL-7B. Runtimes are reported in seconds ($\pm$ standard deviation).
Task Base SC
Mem-Cat-R 1.81 $\pm$ 1.46 5.55 $\pm$ 2.13
Mem-Cat-C 2.06 $\pm$ 1.86 5.37 $\pm$ 1.59
Mem-Loc-R 1.60 $\pm$ 1.20 5.27 $\pm$ 1.99
Mem-Loc-C 2.20 $\pm$ 1.72 5.40 $\pm$ 1.29
Perc-Cat-R 1.54 $\pm$ 1.16 3.21 $\pm$ 0.59
Perc-Cat-C 1.96 $\pm$ 1.43 4.61 $\pm$ 0.60
Perc-Loc-R 1.54 $\pm$ 0.90 2.66 $\pm$ 1.04
Perc-Loc-C 2.07 $\pm$ 1.13 3.69 $\pm$ 0.39
CVR-Cat-H 6.58 $\pm$ 4.67 14.66 $\pm$ 2.86
CVR-Loc-H 6.33 $\pm$ 5.12 14.62 $\pm$ 2.60
CVR-Cat-M 4.76 $\pm$ 4.05 12.59 $\pm$ 1.94
CVR-Loc-M 6.70 $\pm$ 4.12 13.06 $\pm$ 2.10
CVR-Cat-L 3.58 $\pm$ 3.29 10.54 $\pm$ 2.17
CVR-Loc-L 2.83 $\pm$ 3.12 10.76 $\pm$ 1.55
Att-Feat-R 2.74 $\pm$ 1.72 4.06 $\pm$ 0.89
Att-Feat-C 2.91 $\pm$ 1.59 6.09 $\pm$ 1.80
Att-Spa-R 3.06 $\pm$ 1.90 4.66 $\pm$ 1.27
Att-Spa-C 2.78 $\pm$ 1.52 5.89 $\pm$ 1.46
Mem-Dis-Cat-R 3.31 $\pm$ 2.49 6.96 $\pm$ 2.55
Mem-Dis-Cat-C 3.21 $\pm$ 3.18 9.12 $\pm$ 2.89
Mem-Dis-Loc-R 2.77 $\pm$ 2.45 6.51 $\pm$ 2.45
Mem-Dis-Loc-C 3.28 $\pm$ 2.92 8.35 $\pm$ 2.25
A.11 Prompt & Script Examples
"messages": [
"role": "user",
"content": "In this task, we will show you a series of frame images. Each frame will
either be blank (delay frame) or contain one or more 3D objects. The objects will
always be from one of eight categories: benches, boats, cars, chairs, couches,
lighting, planes, and tables. For each category, there are eight unique objects that
could be used in the task. Any object sampled will be displayed as an image taken
from a random viewing angle. The objects will be placed in one of four locations:
top left, top right, bottom left, and bottom right. If there are multiple objects
on a single frame, only one of them would be specified in the task instruction
by either its location or its category. A written instruction will be provided.
Your goal is to follow the instructions and answer the question contained in the
instructions. Answers will always be one of the following: true, false, bottom right,
bottom left, top left, top right, benches, boats, cars, chairs, couches, lighting,
planes, tables.
@@ A.11 Prompt & Script Examples
Task instruction: observe object 1 in frame 1, observe object 2 in frame 2,
observe object 3 in frame 3, observe object 4 in frame 4, observe object 5
in frame 5, delay, observe object 6 in frame 7, delay, observe object 7 in frame 9,
if identity of object 3 equals identity of object 2, then location of object 7
not equals location of object 6 and identity of object 5 equals identity of object 4?
else location of object 1?
[…]
<image><image>
[…]
Provide your answer here: "
"role": "assistant",
"content": "top right"
"images": [
"../trial0/frames/epoch0.png",
"../trial0/frames/epoch0.png",
"../trial0/frames/epoch0.png",
"../trial0/frames/epoch0.png",
"../trial0/frames/epoch0.png",
"../trial0/frames/epoch0.png",
"../trial0/frames/epoch0.png",
"../trial0/frames/epoch0.png",
"../trial0/frames/epoch0.png"
@@
"Frame 1: A chairs located at the top left",
"Frame 2: A chairs located at the top right",
"Frame 3: A benches located at the top left",
"Frame 4: A benches located at the bottom right",
"Frame 5: A boats located at the bottom left",
"Frame 6: A benches located at the bottom right",
"Frame 7: delay frame",
"Frame 8: delay frame",
"Frame 9: A planes located at the top left"
@@
def caption_image(image_path, model, config):
# Prepare a prompt asking the model to caption the image:
caption_prompt = [
"role": "user",
"content":[
"type": "text",
"text": "Please provide a concise caption for the given image, \ including what the location of each the object in the images \ are and what the category of each object is. Each image either \ is blank (a delay frame) or contains one or more 3D objects from \ one of eight categories: benches, boats, cars, chairs, couches, \ lighting, planes, and tables. The object is placed in one of four \ locations: top left, top right, bottom left, or bottom right."
"type": "image_url",
"image_url": {"url": f"data:image/png;base64,{encode_image(image_path)}"}
[…]
response = model.chat.completions.create(
model=config.get("oai-model", "gpt-4o-mini"),
messages=caption_prompt,
max_tokens=1024
[…]
caption = response.choices[0].message.content.strip()
return caption
@@
def evaluate_model(task_base, trial_num, model, config):
trial_base = f"{task_base}/trial{trial_num}/frames/"
task_info_path = os.path.join(trial_base, "new_task_info.json")
[…]
data = json.load(file)
[…]
instruction = data[’new_instruction’]
answers = data[’answers’]
[…]
pattern = epoch*.png
matching_files = [f for f in os.listdir(trial_base) if fnmatch.fnmatch(f, pattern)]
matching_files.sort()
[…]
# NEW PART: First, caption each image
captions = []
for image_path in images:
caption = caption_image(image_path, model, config)
captions.append(caption)
[…]
# Incorporate these captions into the final prompt
# Instead of sending images to the model in the final task, we send their captions:
[…]
# The prompt should now describe that we have textual descriptions of each frame:
possible_answers = reduce(lambda x, y: f"{x}, {y}", get_all_possible_ans(answers[-1]))
[…]
f"""In this task, we will show you a series of frames described by captions instead of direct images. Each frame either is blank (a delay frame) or contains a 3D object from one of eight categories: benches, boats, cars, chairs, couches, lighting, planes, and tables. The object is placed in one of four locations: top left, top right, bottom left, or bottom right.
[…]
@@
messages = [
"role": "user",
"content": [
"type": "text",
"text": prompts[0],
"type": "text",
"text": prompts[1],
[…]
instruction = instruction.join([f"Frame {i+1}: {cap}" for i, cap in enumerate(captions)])
[…]
response = model.chat.completions.create(
model=config.get("oai-model", "gpt-4o-mini"),
messages=messages,
max_tokens=config.get("max_new_tokens", 50),
[…]